OpenStudy (ashking1):

Find the vertex of the graph of the function. f(x) = 3x2 - 24x + 51 can someone help me understand vertex and how to solve for it please

OpenStudy (welshfella):

f(x) = 3(x^2 - 8x + 17) = 3[(x - 4)^2 - 16 + 17] = 3(x - 4)^2 + 3 this is vertex form

OpenStudy (welshfella):

the vertex is at (4,3)

OpenStudy (welshfella):

general form of vertex is y = a(x - b)^2 + c where (a,c) is the coordinates of the vertex of the parabola

OpenStudy (welshfella):